# Cybersecurity of Quantum Key Distribution Implementations

### v2

@@ 5.2 The Reversed-Space Attack
5.2.1 Bob’s Measurement
@@
In order to define the reversed space relevant to this implementation, we start with Bob’s possible outcomes, and we use time reversal to find the exact Hilbert space $\mathcal{H}^{\mathrm{P}}$ which is controlled by Eve and affects Bob’s measurement outcome. First, let $\mathcal{H}^{\mathrm{B}}$ be the span of $\{|{j}\rangle_{\mathrm{B}}\}$, for all basis states $|{j}\rangle_{\mathrm{B}}$ measured by Bob. Then, $\mathcal{H}^{\mathrm{P}}$ is defined to be the span of $\{U_{\mathrm{B}_{s}}^{\dagger}|{j}\rangle_{\mathrm{B}}\}$, for any $s\in\{1,\ldots,m\}$ and all basis states $|{j}\rangle_{\mathrm{B}}\in\mathcal{H}^{\mathrm{B}}$, after tracing out any ancillary space not available to Eve (resulted from an ancillary space $\mathcal{H}^{\mathrm{anc}}$ added by Bob in the “forward in time” description). Any state in Eve’s space that is orthogonal to $\mathcal{H}^{\mathrm{P}}$ goes, after $U_{\mathrm{B}_{s}}$, to a state which is orthogonal to $\mathcal{H}^{\mathrm{B}}$ and can never affect Bob.
Since in the ideal-Alice case, her Hilbert space $\mathcal{H}^{\mathrm{A}}$ is a subspace of $\mathcal{H}^{\mathrm{P}}$, we treat Alice’s qubit as the span of two orthonormal states in $\mathcal{H}^{\mathrm{P}}$.
@@ 5.2.2 Eve’s Attack
$|{0}\rangle_{\mathrm{E}}|{\psi}\rangle_{\mathrm{P}}=\sum_{i}\alpha_{i}|{0}\rangle_{\mathrm{E}}|{i}\rangle_{\mathrm{P}}\stackrel{{\scriptstyle U_{\mathrm{E}}}}{{\longrightarrow}}\sum_{i,k}\alpha_{i}\epsilon_{i,k}|{E_{i,k}}\rangle_{\mathrm{E}}|{k}\rangle_{\mathrm{P}}.$ (6)
@@
$|{k}\rangle_{\mathrm{P}}|{0}\rangle_{\mathrm{anc}}\stackrel{{\scriptstyle U_{\mathrm{B}_{s}}}}{{\longrightarrow}}\sum_{j}\beta^{s}_{k,j}|{j}\rangle_{P\otimes\mathrm{anc}},$ (7)
@@
$|{\Psi_{\mathrm{EB}}}\rangle\triangleq\sum_{i,k,j}\alpha_{i}\epsilon_{i,k}\beta^{s}_{k,j}|{E_{i,k}}\rangle_{\mathrm{E}}|{j}\rangle_{P\otimes\mathrm{anc}},$ (8)
@@
5.2.3 Bob’s Interpretation and Oblivious Attacks
@@
  1. 1.
    Outcomes interpreted as a loss: transmissions that can naturally occur but supply no information to Bob. For example, vacuum pulses that make no detector click, or measurement results that are inconclusive and are part of the theoretical protocol (as occurs in the B92 protocol [2]). These outcomes are denoted as the set $J_{\operatorname{loss}}$.
  2. 2.
@@
It is Bob’s choice of interpretation that determines whether a specific outcome is considered a loss or an invalid result. Generally speaking, when an invalid outcome increases Bob’s measured error rate, we put it in the set $J_{\operatorname{invalid}}$, and when it is ignored by Bob, we put it in $J_{\operatorname{loss}}$. As an example, let us consider the case where Alice’s state generation is not totally ideal, and she can either send a single photon or, at most, two photons. If Bob treats these cases of noticing many photons as a loss (i.e., he ignores that transmission, thus this measurement is in $J_{\operatorname{loss}}$) rather than as an error, this results in a major security hole [55, 60].
@@
$\sum_{i,k}\alpha_{i}\epsilon_{i,k}\beta^{s}_{k,j}|{E_{i,k}}\rangle_{\mathrm{E}}=0,$ (9)
@@
To clarify the notations for oblivious attacks, let us provide a simple example and show that a CNOT attack made by Eve does not satisfy the conditions of Observation 1 and thus can be noticed by Bob. For instance, consider a standard BB84 scheme [1] in which Bob’s setup for the computational basis is $U_{\mathrm{B}_{C}}=I$ the identity, and for the Hadamard basis, $U_{\mathrm{B}_{H}}=H$ is Hadamard transformation; both are followed by a measurement in the computational basis. Assume Alice sends $|{\psi}\rangle_{\mathrm{A}}=|{+}\rangle$, but Eve performs a CNOT attack using the computational basis. After the attack, the system (Alice’s qubit and Eve’s added ancilla) is in the state $|{\tilde{\psi}}\rangle=(|{E_{0,0}}\rangle_{\mathrm{E}}|{0}\rangle_{\mathrm{P}}+|{E_{1,1}}\rangle_{\mathrm{E}}|{1}\rangle_{\mathrm{P}})/\sqrt{2}$ with orthogonal $|{E_{0,0}}\rangle_{\mathrm{E}}$ and $|{E_{1,1}}\rangle_{\mathrm{E}}$. Assume Bob sets his apparatus to the Hadamard basis (same as Alice). Thus, he applies the Hadamard transformation, and measures subsystem $\mathrm{P}$ of the resulting state $|{\Psi_{\mathrm{EB}}}\rangle=(I_{\mathrm{E}}\otimes U_{\mathrm{B}_{H}})|{\tilde{\psi}}\rangle=(|{E_{0,0}}\rangle_{\mathrm{E}}|{+}\rangle_{\mathrm{P}}+|{E_{1,1}}\rangle_{\mathrm{E}}|{-}\rangle_{\mathrm{P}})/\sqrt{2}$ in the computation basis. We now show that Bob has a positive probability of measuring $|{1}\rangle$ (that is, $j=1$), while this outcome is in $J_{\operatorname{error}}$ and indicates an error: using the formulation of Observation 1, Alice’s qubit is given by $\alpha_{0}=\alpha_{1}=1/\sqrt{2}$, Eve’s attack by $\epsilon_{0,0}=\epsilon_{1,1}=1$ and $\epsilon_{0,1}=\epsilon_{1,0}=0$ and Bob’s setup by $\beta^{H}_{0,0}=\beta^{H}_{0,1}=\beta^{H}_{1,0}=1/\sqrt{2}$ and $\beta^{H}_{1,1}=-1/\sqrt{2}$. Indeed, Eq. (9) for $j=1$ gives $\sum_{i,k\in\{0,1\}}\alpha_{i}\epsilon_{i,k}\beta^{H}_{k,1}|{E_{i,k}}\rangle_{\mathrm{E}}=\frac{1}{\sqrt{2}}\cdot 1\cdot\frac{1}{\sqrt{2}}|{E_{0,0}}\rangle+\frac{1}{\sqrt{2}}\cdot 1\cdot\frac{-1}{\sqrt{2}}|{E_{1,1}}\rangle$, which is non-zero since $|{E_{0,0}}\rangle$ and $|{E_{1,1}}\rangle$ are orthogonal.
@@
5.3 Example: a Complete Reversed-Space Attack
A detailed example of a Reversed-Space Attack on a time-based interferometric BB84 scheme, which yields an oblivious attack on that implementation, is given in Appendix B. This subsection gives a short overview of the attack.
In the attacked implementation, Alice encodes her qubit using time-bin encoding with two time-bins (denoted $t^{\prime}_{0}$ and $t^{\prime}_{1}$). However, Bob’s implementation measures six different modes in order to perform his measurement: three different times in the two arms of an interferometer. Furthermore, Reversed-Space analysis reveals that Bob’s measurement can be affected by two additional times, denoted $t^{\prime}_{-1}$ and $t^{\prime}_{2}$.
When analyzing Bob’s interpretation of the two additional times, it is revealed that pulses in the two additional times force Bob to measure in the basis of Eve’s choice (or experience a loss of the signal).
Finally, Eve uses these additional states and their interpretation to construct an attack that has zero chance of generating an erroneous measurement in Bob’s system.
This attack breaks the security (and the robustness: see [61]) of the implementation, because Eve gains full information on the key without inducing any error and without causing the protocol to abort.
Another successful reversed-space attack on a polarization-based interferometric BB84 implementation has been reported in [24].
5.4 Discussion: Countermeasures to Reversed-Space Attacks
@@
5.5 Other Examples of Reversed-Space Attacks
In Appendix D.1 we analyze the setting described in Example 1 of Section 3.2.1 using the Reversed-Space method. This setting is known to be exposed to attacks that are, in fact, a special case of Reversed-Space Attacks.
In Appendix D.2 we revisit the setting of Section 5.3 where Bob measures six modes in the interferometric setup. Now, we restrict Bob to measuring only two modes: the middle time-bin of each interferometer arm. We design an attack that is much stronger than the attack presented in Section 5.3. Its existence strengthens the conclusions presented in Section 5.4, which imply that modifying the measurement scheme of a QKD implementation needs to be thoroughly analyzed.
5.6 Summary: Reversed-Space is an Exploit
@@
Interpretation Vulnerabilities, by Definition 5, reveal additional information on how Bob interprets his incoming states; specifically, they reveal interpretations that should not exist in an ideal protocol, and are necessary information for Eq. (9) that defines the “zero noise requirement”. These interpretations are often the key to the success of the attack.
@@
6 Quantum Side-Channel Attacks
In this section, we examine side-channel attacks, which exploit unintended physical interfaces to compromise security, in the context of QKD devices. As classical intuition fails to capture the nature of side channels in QKD devices, we introduce a new definition of “Quantum Side-Channel Attacks”, and their counterpart which we name “Quantum State-Channel Attacks”. Through concrete examples, we show that Quantum Side-Channel Attacks can target a wide range of device components. Because these attacks are often overlooked in standard security analyses, our definition emphasizes the need to treat them explicitly, providing system designers with a clearer understanding and a stronger basis for defending against them.
6.1 The Classical Concept of Side-Channel Attacks
@@
The most common examples of these interfaces are the physical aspects of the system implementation, such as execution time, electromagnetic radiation, and power consumption. For example, by measuring the exact duration of a cryptographic computation, an attacker can deduce information about the secret key used in the computation [62, 63]. In another example, electromagnetic radiation was applied to a tamper-resistant device to trigger faults in its computation, modifying its behavior and revealing information on the device’s secret key [64, 65].
In fact, side-channel attacks that use physical imperfections are so common that some consider an attack to be a side-channel attack if and only if it is based on the physical implementation of the device [66].
6.2 Quantum Side-Channel Attacks and Quantum State-Channel Attacks
@@
Previous works have used intuition from classical cybersecurity and referred to all attacks on QKD systems that utilize physical imperfections as “side-channel attacks” (see, e.g., [55, 67]). However, there is a clear issue with this approach: it classifies almost every attack as a “side-channel attack”, rendering the definition useless. In QKD implementations, the device that prepares the signals and the device that measures them perform a physical process, and the signals themselves are physical, photonic pulses. Thus, if the criterion for a side-channel attack is to utilize physical imperfections, any attack that utilizes an imperfection in the preparation device, the measurement device, or the signals is a side-channel attack.
@@
In a transmitting device (Alice), the conventional interface is the device’s transmissions: the space of states that the device can transmit. In a measuring device (Bob), the conventional interface is the device’s measurement of incoming signals, which is modeled as the space of states that affect Bob’s measurements. Thus, we view side-channel attacks on QKD as attacks that depend on more than Alice’s sent states, or affect something other than the space that affects Bob’s measurements (or the ancillas Eve can save inside her system). We arrive at the following definition:
@@
  1. 1.
    its input depends on more than Alice’s sent states;
  2. 2.
    its output affects more than the space that affects Bob’s measurements and the attacker’s private ancillas.
@@
  1. 1.
    its input depends only on Alice’s sent states;
  2. 2.
    its output affects only the space that affects Bob’s measurements and the attacker’s private ancillas.
Note that Quantum Side-Channel Attacks can utilize resources that cannot be used in Quantum State-Channel Attacks. In addition, there can be attacks that are neither Quantum Side-Channel Attacks nor Quantum State-Channel Attacks: for example, an attack scenario where Eve installs a camera in Bob’s lab and sees him entering his basis choices (which is typically disallowed by the assumptions of quantum cryptography).
@@
A full, formal definition of the two terms, based on the Hilbert spaces that each device uses, can be found in Appendix E.
There have been past attempts to classify attacks on QKD implementations that do not rely on the conventional interfaces of the devices. Vakhitov, Makarov, and Hjelme [17] used the term “conventional optical eavesdropping” to describe attacks where “Eve can get information by using loopholes in Alice’s and Bob’s optical set-up rather than by measuring the transmitted quantum states”. This term is, in fact, a special case of our definition. While forcing leakage from the optical setups of Alice and Bob certainly qualifies as a side channel, our definition does not require that we not measure the transmitted quantum states. In fact, as shown below, several Quantum Side-Channel Attacks collect data from QKD devices through side channels, or manipulate QKD devices through side channels, in order to perform more successful measurements of the transmitted quantum states. Furthermore, there are other elements besides the optical setup that can be used to launch a Quantum Side-Channel Attack, as explored below.
6.3 Examples

### v3

@@ 5.2. The Reversed-Space Attack
In this subsection, we formally write out the components of the Reversed-Space Attack.
5.2.1. Bob’s Measurement
@@
In order to define the reversed space relevant to this implementation, we start with Bob’s possible outcomes, and we use time reversal to find the exact Hilbert space $\mathcal{H}^{\mathrm{P}}$ which is controlled by Eve and affects Bob’s measurement outcome. First, let $\mathcal{H}^{\mathrm{B}}$ be the span of $\{|{j}\rangle_{\mathrm{B}}\}$, for all basis states $|{j}\rangle_{\mathrm{B}}$ measured by Bob. Then, $\mathcal{H}^{\mathrm{P}}$ is defined to be the span of $\{U_{\mathrm{B}_{s}}^{\dagger}|{j}\rangle_{\mathrm{B}}\}$, for all $s\in\{1,\ldots,m\}$ and all basis states $|{j}\rangle_{\mathrm{B}}\in\mathcal{H}^{\mathrm{B}}$, after tracing out any ancillary space not available to Eve (resulted from an ancillary space $\mathcal{H}^{\mathrm{anc}}$ added by Bob in the “forward in time” description). Any state in Eve’s space that is orthogonal to $\mathcal{H}^{\mathrm{P}}$ goes, after $U_{\mathrm{B}_{s}}$, to a state which is orthogonal to $\mathcal{H}^{\mathrm{B}}$ and can never affect Bob.
Since in the ideal-Alice case, Alice’s Hilbert space $\mathcal{H}^{\mathrm{A}}$ is a subspace of $\mathcal{H}^{\mathrm{P}}$, we treat Alice’s qubit as the span of two orthonormal states in $\mathcal{H}^{\mathrm{P}}$.
@@ 5.2.2. Eve’s Attack
(6) $|{0}\rangle_{\mathrm{E}}|{\psi}\rangle_{\mathrm{P}}=\sum_{i}\alpha_{i}|{0}\rangle_{\mathrm{E}}|{i}\rangle_{\mathrm{P}}\stackrel{{\scriptstyle U_{\mathrm{E}}}}{{\longrightarrow}}\sum_{i,k}\alpha_{i}\epsilon_{i,k}|{E_{i,k}}\rangle_{\mathrm{E}}|{k}\rangle_{\mathrm{P}}.$
@@
(7) $|{k}\rangle_{\mathrm{P}}|{0}\rangle_{\mathrm{anc}}\stackrel{{\scriptstyle U_{\mathrm{B}_{s}}}}{{\longrightarrow}}\sum_{j}\beta^{s}_{k,j}|{j}\rangle_{P\otimes\mathrm{anc}},$
@@
(8) $|{\Psi_{\mathrm{EB}}}\rangle\triangleq\sum_{i,k,j}\alpha_{i}\epsilon_{i,k}\beta^{s}_{k,j}|{E_{i,k}}\rangle_{\mathrm{E}}|{j}\rangle_{P\otimes\mathrm{anc}},$
@@
5.2.3. Bob’s Interpretation and Oblivious Attacks
@@
  1. (1)
    Outcomes interpreted as a loss: transmissions that can naturally occur but supply no information to Bob. For example, vacuum pulses that make no detector click, or measurement results that are inconclusive and are part of the theoretical protocol (as occurs in the B92 protocol (B92, )). These outcomes are denoted as the set $J_{\operatorname{loss}}$.
  2. (2)
@@
It is Bob’s choice of interpretation that determines whether a specific outcome is considered a loss or an invalid result. Generally speaking, when an invalid outcome increases Bob’s measured error rate, we put it in the set $J_{\operatorname{invalid}}$, and when it is ignored by Bob, we put it in $J_{\operatorname{loss}}$. As an example, let us consider the case where Alice’s state generation is not totally ideal, and she can either send a single photon or, at most, two photons. If Bob treats these cases of noticing many photons as a loss (i.e., he ignores that transmission, thus this measurement is in $J_{\operatorname{loss}}$) rather than as an error, this results in a major security hole (gottesman_security_2004, ; hwang_no-clicking_2008, ).
@@
(9) $\sum_{i,k}\alpha_{i}\epsilon_{i,k}\beta^{s}_{k,j}|{E_{i,k}}\rangle_{\mathrm{E}}=0,$
@@
To clarify the notations for oblivious attacks, let us provide a simple example and show that a CNOT attack made by Eve does not satisfy the conditions of Observation 1 and thus can be noticed by Bob. For instance, consider a standard BB84 scheme (BB84, ) in which Bob’s setup for the computational basis is $U_{\mathrm{B}_{\mathrm{C}}}=I$ the identity, and for the Hadamard basis, $U_{\mathrm{B}_{\mathrm{H}}}=H$ is Hadamard transformation; both are followed by a measurement in the computational basis. Assume Alice sends $|{\psi}\rangle_{\mathrm{A}}=|{+}\rangle$, but Eve performs a CNOT attack using the computational basis. After the attack, the system (Alice’s qubit and Eve’s added ancilla) is in the state $|{\tilde{\psi}}\rangle=(|{E_{0,0}}\rangle_{\mathrm{E}}|{0}\rangle_{\mathrm{P}}+|{E_{1,1}}\rangle_{\mathrm{E}}|{1}\rangle_{\mathrm{P}})/\sqrt{2}$ with orthogonal $|{E_{0,0}}\rangle_{\mathrm{E}}$ and $|{E_{1,1}}\rangle_{\mathrm{E}}$. Assume Bob sets his apparatus to the Hadamard basis (same as Alice). Thus, he applies the Hadamard transformation, and measures subsystem $\mathrm{P}$ of the resulting state $|{\Psi_{\mathrm{EB}}}\rangle=(I_{\mathrm{E}}\otimes U_{\mathrm{B}_{\mathrm{H}}})|{\tilde{\psi}}\rangle=(|{E_{0,0}}\rangle_{\mathrm{E}}|{+}\rangle_{\mathrm{P}}+|{E_{1,1}}\rangle_{\mathrm{E}}|{-}\rangle_{\mathrm{P}})/\sqrt{2}$ in the computation basis. We now show that Bob has a positive probability of measuring $|{1}\rangle$ (that is, $j=1$), while this outcome is in $J_{\operatorname{error}}$ and indicates an error: using the formulation of Observation 1, Alice’s qubit is given by $\alpha_{0}=\alpha_{1}=1/\sqrt{2}$, Eve’s attack by $\epsilon_{0,0}=\epsilon_{1,1}=1$ and $\epsilon_{0,1}=\epsilon_{1,0}=0$ and Bob’s setup by $\beta^{\mathrm{H}}_{0,0}=\beta^{\mathrm{H}}_{0,1}=\beta^{\mathrm{H}}_{1,0}=1/\sqrt{2}$ and $\beta^{\mathrm{H}}_{1,1}=-1/\sqrt{2}$. Indeed, Eq. (9) for $j=1$ gives $\sum_{i,k\in\{0,1\}}\alpha_{i}\epsilon_{i,k}\beta^{\mathrm{H}}_{k,1}|{E_{i,k}}\rangle_{\mathrm{E}}=\frac{1}{\sqrt{2}}\cdot 1\cdot\frac{1}{\sqrt{2}}|{E_{0,0}}\rangle+\frac{1}{\sqrt{2}}\cdot 1\cdot\frac{-1}{\sqrt{2}}|{E_{1,1}}\rangle$, which is non-zero since $|{E_{0,0}}\rangle$ and $|{E_{1,1}}\rangle$ are orthogonal.
@@
5.3. Example: a Complete Reversed-Space Attack
Appendix B provides a detailed description of a time-based interferometric BB84 implementation, which appears in (Walton2003, ; Nambu2004, ; Jaeger2006, ; Sasaki2011, ), and constructs a successful Reversed-Space Attack on that implementation. This subsection gives a short overview of the implementation and the attack.
In the attacked implementation, Alice encodes her qubit using time-bin encoding with two time-bins (denoted $t^{\prime}_{0}$ and $t^{\prime}_{1}$). Bob’s interferometer-based device splits Alice’s pulse into six separate modes: two output arms of the interferometer (denoted $s$ and $d$), where in each arm, a photon could appear in three possible time-bins (denoted 0, 11, and $2$). Thus, Bob’s detectors measure six different modes which can be affected by Alice’s qubit.
A reversed-space analysis reveals that Bob’s measurement is not only affected by Alice’s qubit, but also by two additional times at the interferometer’s input, denoted $t^{\prime}_{-1}$ and $t^{\prime}_{2}$. The analysis of Bob’s operation on these two additional times reveals that carefully-crafted pulses in the two additional times can force Bob to either measure in a basis of Eve’s choice, or experience a loss of the signal. Eve can use these additional modes to design an attack that never causes an erroneous measurement in Bob’s system.
This attack breaks the security (and the robustness: see (BKM2007, )) of this BB84 implementation, because Eve gains full information on the key without inducing any error and without causing the protocol to abort.
We note that if Bob is constrained to only measure two specific modes instead of all the six resulting modes, a simpler and stronger attack can be devised; see (gelles_reversed_2016, , Section III D) for more details. We also note that another successful reversed-space attack on a polarization-based interferometric BB84 implementation was reported in (boyer_fixedapp_2014, ).
5.4. Discussion: Countermeasures to Reversed-Space Attacks
@@
5.5. Summary: Reversed-Space is an Exploit
@@
Interpretation Vulnerabilities, by Definition 5, reveal additional information on how Bob interprets his incoming states; specifically, they reveal interpretations that should not exist in an ideal protocol, and are necessary information for Eq. (9) that defines the “zero noise requirement” of Observation 1. These interpretations are often the key to the success of the attack.
@@
6. Quantum Side-Channel Attacks
In this section, we examine and define the notion of “side-channel attacks” in the context of QKD devices. Similarly to the classical side-channel attacks, these attacks exploit unintended physical interfaces that potentially compromise the security of the underlying QKD protocol. However, unlike classical side-channel attacks, physical interfaces of QKD protocols are often part of the acceptable communication between Alice and Bob.
Therefore, we introduce new definitions of “Quantum Side-Channel Attacks” and their complement, which we name “Quantum State-Channel Attacks”. We also present concrete examples to show that Quantum Side-Channel Attacks can target a wide range of device components. Because these attacks are often overlooked in standard security analyses, our definition emphasizes the need to treat them explicitly and provides system designers with a clearer understanding and a stronger basis for defending against them.
6.1. The Classical Concept of Side-Channel Attacks
@@
The most common examples of these interfaces are the physical aspects of the system implementation, such as execution time, electromagnetic radiation, and power consumption. For example, by measuring the exact duration of a cryptographic computation, an attacker can deduce information about the secret key used in the computation (Kocher96, ; Brumley11, ). In another example, electromagnetic radiation was applied to a tamper-resistant device to trigger faults in its computation, modifying its behavior and revealing information on the device’s secret key (Boneh97, ; Biham97, ).
In fact, side-channel attacks that use physical imperfections are so common that some consider an attack to be a side-channel attack if and only if it is based on the physical implementation of the device (NIST-SP-800-63-3, ).
6.2. Quantum Side-Channel Attacks and Quantum State-Channel Attacks
@@
Previous works have used intuition from classical cybersecurity and referred to all attacks on QKD systems that utilize physical imperfections as “side-channel attacks” (see, e.g., (Nauerth2009, )). However, there is a clear issue with this approach: it classifies almost every attack as a “side-channel attack”, rendering the definition useless. In QKD implementations, the device that prepares the signals and the device that measures them perform a physical process, and the signals themselves are physical, photonic pulses. Thus, if the criterion for a side-channel attack is to utilize physical imperfections, any attack that utilizes an imperfection in the preparation device, the measurement device, or the signals is a side-channel attack.
@@
In a transmitting device (Alice), the conventional interface is the device’s transmissions: the space of states that the device can transmit. In a measuring device (Bob), the conventional interface is the device’s measurement of incoming signals, which is modeled as the space of states that affect Bob’s measurements. When Eve performs an attack, the attack’s input includes the states that arrive from Alice’s device, the implementation’s physical environment, and Eve’s private ancillas, while the attack’s output includes the states it sends to Bob’s device, the physical environment, and Eve’s private ancillas.
Thus, we view side-channel attacks on QKD as attacks that depend on more than Alice’s sent states, or affect something other than the space that affects Bob’s measurements (or the ancillas Eve can save inside her system). We arrive at the following definition:
@@
  1. (1)
    its input non-trivially depends on more than the states Alice sends to Bob,
  2. (2)
    or its output non-trivially affects more than the space that affects Bob’s measurements and the attacker’s private ancillas.
@@
  1. (1)
    its input non-trivially depends only on the states Alice sends to Bob,
  2. (2)
    and its output non-trivially affects only the space that affects Bob’s measurements and the attacker’s private ancillas.
Note that Quantum Side-Channel Attacks can utilize resources that cannot be used in Quantum State-Channel Attacks. However, those resources are limited by standard working assumptions on the implementation, as described in international standards such as (iso-qkd-standard-pt-1, ). Specifically, when considering attacks in this paper, we assume that each QKD device operates inside a shielded environment (while the channels connecting them are not shielded); that the users of the legitimate devices (Alice and Bob) are trusted; and that the pre-shared secrets of the devices are secure. Under these assumptions, the definitions of Quantum Side-Channel Attacks and Quantum State-Channel Attacks are complementary. However, there can exist attacks that violate these assumptions: for example, an attack scenario where Eve installs a camera in Bob’s lab and sees him entering his basis choices.
@@
Further note that Definitions 1 and 2 are somewhat informal: a rigorous formulation would require a precise mathematical definition of Alice’s and Bob’s enlarged Hilbert spaces corresponding to the physical reality where both Alice and Bob could be imperfect, in addition to Eve’s most general attack on them. The precise definition of those spaces and their relations is complex and subtle (see, e.g., (gelles_qsa_2007, ) for such an attempt) and is left for future work.
There have been past attempts to classify attacks that do not rely on the conventional interfaces of the QKD devices. Vakhitov, Makarov, and Hjelme (vakhitov_large_2001, ) used the term “conventional optical eavesdropping” to describe attacks where “Eve can get information by using loopholes in Alice’s and Bob’s optical set-up rather than by measuring the transmitted quantum states”. This term is, in fact, a special case of our definition. While forcing leakage from the optical setups of Alice and Bob certainly qualifies as a side channel, our definition does not require that we not measure the transmitted quantum states. In fact, as shown below, several Quantum Side-Channel Attacks collect data from QKD devices through side channels, or manipulate QKD devices through side channels, in order to perform more successful measurements of the transmitted quantum states. Furthermore, there are other elements besides the optical setup that can be used to launch a Quantum Side-Channel Attack, as explored below.
6.3. Examples